OpenStudy (anonymous):

Need help!!! Please Describe how to prepare a 10.0% w/v solution of salt in water in a 100mL volumetric flask.

OpenStudy (cuanchi):

10.0 w/v means 10.0 g of solute in 100 mL of solution. see attachment

OpenStudy (ciarán95):

The weight/volume (w/v) percent measurement for the concentration of a solution can be found as follows: \[\[w/v% = \frac{ Weight of solute (g) }{ Volume of solution (ml) } \times 100\] It is essentially the fraction of our solution which is the solid solute we're trying to dissolve, compared to the amount of solvent present. So, if we have a 100 mL volumetric flask, we can measure out 100 mL of our solvent (i.e. prepare 100 mL of a solution). If we wanted this solution to be 10% w/v: \[\frac{ Weight~of~solute (g) }{ 100 mL } \times 100 = 10\] So, \[Weight~of~solute~(g) = \frac{ (100)(10) }{ 100 }\]

OpenStudy (ciarán95):

Sorry, the equation didn't seem to come out properly for me there: \[\frac{ weight }{volume} = \frac{ Weight~of~solute(g) }{ Volume~of~solvent(mL) }\times100\] This should tell us how much of our solute we need to add to our known amount of solvent, thus enabling us to describe how we would prepare our solution.

OpenStudy (kash_thesmartguy):

@Jaynator495